OpenStudy (anonymous):

parallel to the graph of 3x-5y=7 and passes through (0,-6). Write an equation in slope intercept form of the line satisfying the given conditions.

OpenStudy (anonymous):

I'll give the steps to do this. A. Find the slope. If it is parallel to the other graph, then it must have the SAME slope. I'd change the equation to y=mx+b where m equals the slope and b equals the y intercept. So, whatever the coefficient is in this equation will be the slope. B. Now, we need to make the new equation y=mx+b You've already found the slope (m) fill that in. y=?x+b b is the y intercept... Lucky you, your point IS the y intercept (-6). So that gaves you the b... y equals the slope you found above * x + (-6)

OpenStudy (anonymous):

so it is y=3x+6, right?

OpenStudy (texaschic101):

3x - 5y = 7 -5y = -3x + 7 y = 3/5x - 7/5 your slope is 3/5. A parallel line will have the same slope. now we will use y = mx + b and solve for b y = mx + b slope(m) = 3/5 (0,-6) x = 0 and y = -6 now we sub y = mx + b -6 = 3/5(0) + b -6 = b therefore, your parallel line is : y = 3/5x - 6

OpenStudy (anonymous):

Thank you for showing me the steps that help very much!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (texaschic101):

glad I could help :)

OpenStudy (anonymous):

thank you again